Angie has 12 cookies ana a marginal value curve described by P=45-3Qa,while Benny has 0 cookies and a marginal value curve described by P=44-2Qb.
Assume Angie and Benny are the only two people in the cookie market .
a) Write down the market demand function and market supply function and draw the market demand and supply curves in one graph .
b) Solve for the market equilibrium price and quantity.
c) How many cookies does each of them consume ?
d) How many cookies are traded ?
e) Solve for the consumer surplus and seller surplus .
Answer #1

a) Market demand is sum of all individual demands.

P=45+3Q+(44-2Q)

=45+3Q+44-2Q

b) Market equilibrium and price:-

P=45+3Q

P=44-2Q

Dd=Ss

45+3Q=44-2Q

45-44=-2Q-3Q

1=-5Q

Q=1/-5

=-0.2

P=45+3*-0.2

=45-0.6

=44.4

c) 44-2Q

=44-2*-0.2

=44+0.4

=44.4